OpenStudy (anonymous):

Find the solutions to the equation below. Check all that apply. 30x^2-26x+4=0 A.x=1/2 B.X=1/5 C.X=3/5 D.X=1/3 E.X=2/5 F.X=4/5

OpenStudy (anonymous):

i got B. nd E.

OpenStudy (anonymous):

yes for B :) You totally got this!

OpenStudy (anonymous):

\[30x^2-26x+4=0 \] NOT E

OpenStudy (anonymous):

\[30x^2-26x+4=0 \]\[2 (3 x-2) (5 x-1) = 0\]

OpenStudy (anonymous):

i wrote it wrong its suppose t0 be 2/3

OpenStudy (anonymous):

2/3 AND 1/5

OpenStudy (anonymous):

it suppose to be B.1/5 nd E.2/3

OpenStudy (anonymous):

CORRECT